A hollow, thin-walled cylinder and a solid sphere start from rest and roll without slipping down an inclined plane of length 3.0 m. The cylinder arrives at the bottom of the plane 2.8 s after the sphere. Determine the angle between the inclined plane and the horizontal.
Answer #1

Lets consider a solid rigid body rolling without slipping. The Newtons second law: F = ma mgsino - fr=ma (1) The second New

Then: ts + 2.8 ts J7 Solving for ts ts = 14.34s Using equation (4) 14(3.0) 14.34 = 5* 9.8sino Solving for a O = 0.239°